OpenStudy (anonymous):

Can someone help me with this one problem?(: will fan and medal sqrt 50x^7y^7 * sqrt 6xy^4

OpenStudy (johnweldon1993):

\[\large \sqrt{50x^7y^7} \times \sqrt{6xy^4}\] Like that?

OpenStudy (anonymous):

yes!:)

OpenStudy (johnweldon1993):

Alright, well a couple ways to do this..but to make sure there's no giant number to deal with...lets start off with a quick simpification \[\large \sqrt{50x^7y^7} \times \sqrt{6xy^4}\] Can be written as \[\large \sqrt{(25\times 2) x^7y^7} \times \sqrt{6xy^4}\] All I changed was 50 is now expressed as 25 times 2

OpenStudy (johnweldon1993):

So now...with that 25...we can look at everything separate so we can just focus on the \[\large \sqrt{25}\] Notice anything? can that be simplified?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

5

OpenStudy (johnweldon1993):

Great! And okay...so back to the whole original question \[\large \sqrt{(25\times 2)x^7y^7} \times \sqrt{6xy^4}\] Now that we know the square root of 25 is 5...we can take the '5' out of the square root \[\large 5\sqrt{2x^7y^7} \times \sqrt{6xy^4}\] That make sense?

OpenStudy (anonymous):

yeah!

OpenStudy (johnweldon1993):

Okay cool... Next step, remember that when you multiply square roots...everything under both square roots is now under 1 and being multiplied. Example \(\large \sqrt{2}\times \sqrt{3} = \sqrt{2\times 3}\) okay?

OpenStudy (anonymous):

okay

OpenStudy (johnweldon1993):

Soooo with your question here...we have \[\large 5\sqrt{2x^7y^7} \times \sqrt{6xy^4}\] And when we multiply everything...we have \[\large 5\sqrt{2x^7y^7\times 6xy^4}\] just combined em under the radical

OpenStudy (johnweldon1993):

So now...we just need to multiply everything together... what is \(\large 2x^7y^7 \times 6xy^4\) ?

OpenStudy (anonymous):

i got 12x^8y^11

OpenStudy (johnweldon1993):

Perfect :) Okay...now remember all that is under the radical...so we have \[\large 5\sqrt{12x^8y^{11}}\] Remember how I broke up 50 into 25 times 2 because the 25 has an easy square root? can 12 be broken up in the same way?

OpenStudy (anonymous):

yeah

OpenStudy (johnweldon1993):

And how can that be broken up?

OpenStudy (anonymous):

3 and 4

OpenStudy (anonymous):

wait..i dont know D:

OpenStudy (johnweldon1993):

No no you got it :DD

OpenStudy (anonymous):

oh okay!

OpenStudy (johnweldon1993):

Because 4 has an easy square root of 2 :D

OpenStudy (johnweldon1993):

So we have \[\large 5\sqrt{(4\times 3)x^8y^{11}}\] which we can then write as \[\large (5\times 2)\sqrt{3x^8y^{11}}\] When we take the 2 out of the square root...it then gets multiplied to that 5 we already had out there...so obviously 5 times 2 is 10 so \[\large 10\sqrt{3x^8y^{11}}\]

OpenStudy (johnweldon1993):

Still with me? :)

OpenStudy (anonymous):

yeah!:)

OpenStudy (johnweldon1993):

Great!... Alright...now is where it gets...odd? We can only continue simplifying if we assume that 'x' and 'y' are positive.. But as of now...this is the final simplification.

OpenStudy (johnweldon1993):

IF,and that is only depending on your assignment,we do assume that 'x' and 'y' are positive...then we can just do the last step...so you let me know if you would like to see this part...or if it's not needed :)

OpenStudy (anonymous):

well, it's not really needed but I would still like to see it to get the idea:)

OpenStudy (johnweldon1993):

Okay :) So, assuming that 'x' and 'y' are positive...we have \[\large 10\sqrt{3x^8y^{11}}\] Sidebar: How else can we write \(\large \sqrt{x}\) ? More specifically...in terms of exponents!

OpenStudy (anonymous):

erm, im not sure

OpenStudy (johnweldon1993):

\[\Large \sqrt{x} = x^{\frac{1}{2}}\]

OpenStudy (johnweldon1993):

You can definitely read more about this because it DOES come in handy But for now..take my word for it :D Okay...back to your problem \[\large 10\sqrt{3x^8y^{11}}\] We can write that as \[\Large 10(3x^8y^{11})^\frac{1}{2}\]

OpenStudy (johnweldon1993):

What happens when we have an exponent on the outside of a parenthesis...what do we do? \[\large (x^2)^3\] What is that?

OpenStudy (anonymous):

would it be x^6

OpenStudy (johnweldon1993):

Sorry hun I have to head out for a bit...so I'm gonna do a quick summary So to answer that last thing...we multiply instead of add like normal.. \[\large (x^2)^3 = x^6\] Knowing that...when we look at your question.. \[\large 10(3x^8y^{11})^{\frac{1}{2}}\] We can raise everything in the parenthesis to that one-half power...so \[\large 10(3^{1/2})(x^8)^{1/2}(y^{11})^{1/2}\] What is 3 to the 1/2 power? well it is just the same...3 to the 1/2 power...or (as we now know...is the square root of 3...so we have \[\large 10\sqrt{3}(x^8)^{1/2}(y^{11})^{1/2}\] What is \(\large (x^8)^{1/2}\) ?? we just learned we multiply the 8 and the 1/2 so we have \[\large 10\sqrt{3}x^{4}(y^{11})^{1/2}\] Now finally we just multiply the 11 times the 1/2 and we get \[\large 10\sqrt{3}x^4y^{11/2}\] Which is TRUELY the final simplification! If you still need help on this...feel free to tag another helper...or tag me back and I'll take a look when I return :)

OpenStudy (anonymous):

Well thank you so much for your help! I really appreciate it:)

OpenStudy (johnweldon1993):

Of course! Anytime :)